OpenStudy (anonymous):

I don't get this one. The difference between two numbers is 29. The sum of five times the smaller and two times the larger is 156. Find the two numbers.

OpenStudy (anonymous):

a. 14 and 15 b. -14 and 15 c. -15 and 14 d. 14 and 43

OpenStudy (anonymous):

I don't' get this. :( It doesn't make sense to me all at.

OpenStudy (anonymous):

let a = larger and b = smaller a-b=29; 2a+5b=156 if you know how to use substitution, you should be able to go from there if not, let me know and i'll explain the rest

OpenStudy (anonymous):

okay, give me a second.

OpenStudy (anonymous):

i don't get it. :(

OpenStudy (anonymous):

okay, well first, do you understand what I did there?

OpenStudy (anonymous):

not really. Sorry

OpenStudy (anonymous):

the difference between larger and smaller is 29 means larger - smaller =29

OpenStudy (anonymous):

now for larger just make it a; so a=larger, b=smaller, from now on we use a and b

OpenStudy (anonymous):

so a (larger) - b (smaller) =29

OpenStudy (anonymous):

okay. I get that part.

OpenStudy (anonymous):

the sum means adding so: 5b+2a=156

OpenStudy (anonymous):

i.e. 5(smaller)+2(larger)=156

OpenStudy (anonymous):

okay. Got it.

OpenStudy (anonymous):

so again your two equations are a-b=29 5b+2a=156 so, move b to the right side in equation 1 getting: a=29+b then plug the value (29+b) for a in to the second equation. you should then be able to solve for b, once you get b plug the value for b into either equation to get a. a and b will be your two numbers

OpenStudy (anonymous):

oh okay. i think i can take it from here. Thanks so much!!

OpenStudy (anonymous):

no problem , just following those steps exactly should take you straight to the answer:) as long as they make sense to you

OpenStudy (anonymous):

it does. i found it! :D